An ion having charge +6e is traveling horizontally to the left at 8.15km/s when it enters a magnetic field that is perpendicular to its velocity and deflects it downward with an initial magnetic force of 6.84
Answer #1

A)
The field is directed into the screen.

B)

F = q*v*B*sin(90)

B = F/(q*v)

= 6.84*10^-15/(6*1.6*10^-19*8150)

= 0.874 T
Answer #2

a)directed in to the screen

b) 0.8742 T
Answer #3

Lorrentz equation : F=q(VxB).

B=(6.84*10^15)/(6*1.6*10^(-19)*8150)

B= 8.742X10^(29).

It is directed into the screen.